# Bilateral Propagation Network for Depth Completion

### v1

@@ 1 Introduction
Dense depth perception, i.e. estimating per-pixel distance to the camera, is crucial for 3D tasks, with a rapid growth of applications, such as augmented reality, autonomous driving and robotics. However, due to current hardware limitations, it is still difficult to directly measure dense depth maps. Depth measured from LiDAR [9] or SfM [39] is typically sparse, which is insufficient for real applications like scene reconstruction [48] and robot grasping [30]. Depth completion111 The exact name should be image guided depth completion, if considering some early attempts don’t utilize color images., i.e. estimating dense depth map from sparse depth measurements with a synchronized color image, is a promising solution with the advantage of exploiting complementary information from data in different modalities. A key challenge in this task is to effectively process irregularly sampled sparse depth points.
Refer to caption (a) Refer to caption (b) Refer to caption (c)
@@ 2 Related Work
Depth Refinement. Directly regressed depth map may suffer from blur effect on object boundaries. Depth refinement mostly follows the spatial propagation mechanism [25], which iteratively refine the regressed depth by a local linear model with learned affinity. Cheng et al. [3] adopt convolution operation to update all pixels simultaneously with a fixed kernel size, and further extend it to CSPN++ [4] assembling results from multiple kernel sizes and multiple iteration steps. Park et al. [32] provide a non-local neighborhood for propagation by learning the offsets to the regular grid. Lin et al. [23] extend the propagation process with affinity matrix adaptive to neighbors of different distances. Liu et al. [26] construct graph neural network integrating 3D information to estimate neighbors in each update iteration. Wang et al. [47] adjust the kernel scope from large to small during the update process. These propagation-based methods dominate current SOTA, but still require an initial dense depth which is mostly yielded by directly applying convolutional layers on the sparse depth map. In contrast, our BP-Net propagates depth at the earliest stage to avoid convolving on sparse depth map. Though we implement a simple CSPN++ [4] for depth refinement, our method should be also compatible to other depth refinement approaches.
@@ 3.2.2 Parameter Generation
The $\mathbb{F}$ is conditioned on prior encodings including the image encoding $\mathcal{I}$ at target pixel $i$ and source pixel $j$, the depth encoding $\mathcal{S}$ at source pixel $j$, and the spatial offset encoding $\mathcal{O}$ for pixel coordinates from $i$ to $j$. The spatial offset encoding $\mathcal{O}$ serves as the spatial distance term. While we do not explicitly design the radiometric difference term, we expect that the MLP $\mathbb{F}$, with the ability to approximate a universal function [13], can implicitly consider radiometric difference with all the relative content encodings as input.
@@
The paradigm of content-dependent parameter generation has been used in previous depth completion methods and gained robust results. GuideNet [42] and RigNet [50] leverage content-dependent weights to establish a sophisticated multi-modal feature fusion. CSPN [3] and CSPN++ [4] dynamically generate content-dependent affinity matrix to iteratively enhance the initial depth. There are two notable differences between these existing methods and ours. Firstly, the content-dependent parameters are only learned from image content in these methods, typically neglecting the spatial distance between the target pixel and nearby depth measurements. In contrast, our method takes both image content and spatial distance into account. This allows the propagation of depth with the preference for nearest values in both radiometric and spatial domains. Secondly, the content-dependent parameters are typically employed in the multi-modal fusion or refinement stage in existing methods while we extend this paradigm to the earliest pre-processing stage, enabling subsequent stages to overcome the sparsity problem more effectively.
@@ 3.2.3 Prior Encoding
The proposed bilateral propagation module is arranged in a multi-scale scheme, with prior encodings from the corresponding resolution. The image encoding $\mathcal{I}$ in scale $s$ can be written as:
[…]
Here, for the lowest resolution with $s=5$, we directly adopt the image feature $\mathbf{I}^{s}$ as image encoding. Otherwise, to make image encoding representative, we concatenate the multi-modal fused feature $\mathbf{F}^{s+1}$ (introduced in Sec. 3.3) with depth feature $\mathbf{D}^{s+1}$ in scale $s+1$. The depth feature $\mathbf{D}^{s+1}$ is achieved by inverse projecting estimated depth $D^{s+1}$ to camera space. Then we utilize deconvolution operation $\mathbb{D}$ to upsample the concatenated feature map to scale $s$. Finally, we concatenate the upsampled feature with $\mathbf{I}^{s}$, and adopt an extra convolution operation $\mathbb{C}$ to produce the image encoding.
We inverse project sparse depth map $S^{s}$ to camera space as depth encoding $\mathcal{S}^{s}$. Except $S^{0}=S$, the sparse depth map $S^{s}$ is downsampled via a weighted pooling, whose weights are generated from the image encoding $\mathcal{I}^{s}$. Specially, we adopt periodic shuffling operator to rearrange the feature map from shape ${\frac{H}{s}\times\frac{W}{s}\times s^{2}}$ to shape $H\times W$, and apply exponential transformation, that guarantees the generated weight map $w^{s}$ is positive and has the same resolution as $S$. Then the sparse depth map is downsampled as:
$S_{i}^{s}=\frac{\sum\limits_{j=\mathcal{N}_{s}(i)}w_{j}^{s}S_{j}}{\sum\limits_{j=\mathcal{N}_{s}(i)}w_{j}^{s}\mathbb{I}(S_{j})+\epsilon},$ (4)
where $\mathbb{I}$ is the indicator function for the validity of depth measurement, $\epsilon$ is a value close to 00 to prevent division by zero, and $\mathcal{N}_{s}(i)$ is the set $\{(x,y)|si_{x}\leq x<si_{x}+s,si_{y}\leq y<si_{y}+s\}$, representing pixels with coordinates $(x,y)$ in original resolution corresponding to pixel $i$ in scale $s$, It is worth to note that though $S^{s}$ may be denser in low resolution, it still has invalid pixels when $\sum_{j=\mathcal{N}_{s}(i)}w_{j}S_{j}=0$ in Eq. 4, and only the valid pixels in $S^{s}$ is adopted for depth propagation in Eq. 1 and Eq. 2.
We adopt the directed distance from the target pixel to the valid source pixel, i.e. a 2D vector representing offsets along two axes of the image plane, as the spatial distance encoding $\mathcal{O}$. Position encoding has been used in some recent popular network structures, e.g. Transformer [46] and NeRF [31]. In these networks, Fourier transform or learned embedding is adopted to encode the position to a high dimensional space. From our experiment, we find high dimensional position encoding can induce little performance improvement in our formulation, despite the increased computational cost. It may be because a simple spatial distance term is sufficient for coefficient generation in such a local combination model, similar to bilateral filtering.
@@ 3.3 Multi-Modal Fusion
For each scale, we can get a dense depth map ${{D}^{\prime}}^{s}$ from the bilateral propagation module which is suitable for CNN to process. We adopt image encoding $\mathcal{I}^{s}$ as the image feature, and inverse project ${{D}^{\prime}}^{s}$ to camera space as depth feature. Then, we follow the early-fusion scheme that simply concatenates image feature and depth feature, and feeds the concatenated feature as input to a U-Net structure [38] for multi-modal fusion. The adopted U-Net is an encoder-decoder CNN, aggregating local and global features in multi-scale. The encoder part consists of 2 ResNet blocks [11] for feature extraction in each scale and convolution layer with stride 2 to reduce the resolution of feature map. The decoder part consists of deconvolution layer with stride 2 to upsample feature map, and skip connection to fuse upsampled feature and encoded feature with same resolution by concatenation operation. The U-Net finally yields a fused feature map $\mathbf{F}^{s}$. By applying convolution layer on $\mathbf{F}^{s}$ to estimate the residual depth map, we obtain ${{D}^{\prime\prime}}^{s}={{D}^{\prime}}^{s}+\mathbb{C}(\mathbf{F}^{s})$.
@@ 3.5.2 Training Setting
We implement our method in pytorch [33] and train it on a GPU workstation with 4 Nvidia RTX 3090 GPU. Our network is mainly stacked with ResNet blocks [11], in which a DropPath [21] is added before residual addition as a regularization for training. We adopt AdamW [27] with 0.05 weight decay as the optimizer, and clip gradient whose $l^{2}$-norm is larger than $0.1$. Our method is trained from scratch in roughly $300K$ iterations with OneCycle learning rate policy [41] gradually reducing the learning rate to $25\%$ of the largest learning rate. We set batch size as $8$ and the largest learning rate as $0.001$ for KITTI dataset. For NYUv2 dataset, whose image resolution is lower, we increase the batch size to $16$ and the largest learning rate to $0.002$ correspondingly. The final model is yielded by Exponential Moving Average (EMA) with $0.9999$ decay.
@@ 4.1 Datasets & Evaluation Metrics
Outdoor scene. The KITTI depth completion (DC) dataset [45] is collected by an autonomous driving vehicle, whose ground truth depth is from temporally registered LiDAR scans, further verified by stereo image pairs. This dataset provides $86,898$ and $1,000$ frames for training and validation respectively. Another $1,000$ frames for testing are evaluated on the remote server with a public leaderboard [^2] http://www.cvlibs.net/datasets/kitti/eval_depth.php?benchmark for ranking. We randomly crop frames to $256\times 1216$ for training and directly use the full-resolution frames as input for testing. The standard evaluation metrics are root mean squared error (RMSE), mean absolute error (MAE), root mean squared error of the inverse depth (iRMSE), and mean absolute error of the inverse depth (iMAE), among which RMSE is chosen as the primary metric for ranking.
@@ 4.2 Comparison with SOTAs
We evaluate our method on the test set of the NYUv2 dataset and KITTI DC dataset separately. Tab. 1 lists the quantitative comparison of our method and other top-ranking published methods. On the KITTI DC leaderboard, our method ranks 1st and exceeds all other methods under the primary RMSE metric at the time of paper submission. It also has comparable performance under other evaluation metrics. On the NYUv2 dataset, our method achieves the best RMSE and the best $\delta_{1.25}$.
@@ 4.3.1 Different Number for Neighboring Pixels
In our bilateral propagation module, the target depth is propagated from $N$ nearest neighboring valid depth measurements according to our depth parametrization in Eq. 1. We experiment with different values of $N$ and plot the corresponding RMSEs in Fig. 5. The performance gain is most noticeable when $N$ increases from 1 to 2, after which it gradually decreases and trends toward saturation. Finally, we choose $N=4$ in our method.
4.3.2 Effect of Bilateral Parameters Generation
Table 2: Ablation studies on NYUv2 dataset.
@@ 4.3.3 Effect of Multi-Scale Depth Estimatation
We also verify the effectiveness of our multi-scale scheme, which estimates dense depth in a multi-scale architecture with a multi-scale loss. We train a variant by replacing the multi-scale loss with a loss only on the original scale and train another variant by further replacing the multi-scale architecture with depth estimation only on the original scale. From the results listed in Tab. 2, we can see combining multi-scale architecture and multi-scale loss can improve system performance. Simply introducing the multi-scale architecture without multi-scale loss decreases the performance, which may be caused by the difficulty in training.
@@ 4.3.4 Effect of 3-stage Depth Estimation
At each scale, dense depth is estimated via three sequential stages, consisting of pre-processing (Pre.), multi-modal fusion (MF.), and post-processing (Post.). We train different variants with different combinations of these three stages and analyze the performance gain on RMSE and extra cost on parameters (Params.), Multiply-adds operations (Madds.), GPU RAM and runtime. As shown in Tab. 3, combining MF. with Pre. can achieve $2.12mm$ performance gain on RMSE at the cost of $1.11M$ more Params. and $7.39G$ more Madds. As a comparison, combining MF. with Post. produces less performance gain with more cost on Params. and Madds. In addition, Pre.+MF. achieves better RMSE than MF.+ Post. with less GPU RAM and runtime. This comparison verifies the effectiveness of our 3-stage design and proves the importance of depth propagation at the preprocessing stage in contrast to the refinement stage.
4.4 Input Sparsity Analysis
Table 3: Effect of 3-stage estimation.
@@ 5 Conclusion
Acknowledgements: This work was supported in part by National Natural Science Foundation of China (NSFC) under Grants 62273353 and 62103431.

### v2

@@ 1 Introduction
Dense depth perception, i.e. estimating per-pixel distance to the camera, is crucial for 3D tasks, with a rapid growth of applications, such as augmented reality, autonomous driving and robotics. However, due to current hardware limitations, it is still difficult to directly measure dense depth maps. Depth measured from LiDAR [9] or SfM [38] is typically sparse, which is insufficient for real applications like scene reconstruction [47] and robot grasping [29]. Depth completion [^1] The exact name should be image guided depth completion, if considering some early attempts don’t utilize color images., i.e. estimating dense depth map from sparse depth measurements with a synchronized color image, is a promising solution with the advantage of exploiting complementary information from data in different modalities. A key challenge in this task is to effectively process irregularly sampled sparse depth points.
@@ 2 Related Work
Depth Refinement. Directly regressed depth map may suffer from blur effect on object boundaries. Depth refinement mostly follows the spatial propagation mechanism [24], which iteratively refines the regressed depth by a local linear model with learned affinity. Cheng et al. [3] adopt convolution operation to update all pixels simultaneously with a fixed kernel size, and further extend it to CSPN++ [4] assembling results from multiple kernel sizes and multiple iteration steps. Park et al. [31] provide a non-local neighborhood for propagation by learning the offsets to the regular grid. Lin et al. [22] extend the propagation process with affinity matrix adaptive to neighbors of different distances. Liu et al. [25] construct graph neural network integrating 3D information to estimate neighbors in each update iteration. Wang et al. [46] adjust the kernel scope from large to small during the update process. These propagation-based methods dominate current SOTA, but still require an initial dense depth which is mostly yielded by directly applying convolutional layers on the sparse depth map. In contrast, our BP-Net propagates depth at the earliest stage to avoid convolving on sparse depth map. Though we implement a simple CSPN++ [4] for depth refinement, our method should be also compatible to other depth refinement approaches.
@@ 3.2.2 Parameter Generation
The $\mathbb{F}$ is conditioned on prior encodings including the image encoding $\mathcal{I}$ at target pixel $i$ and source pixel $j$, the depth encoding $\mathcal{S}$ at source pixel $j$, and the spatial offset encoding $\mathcal{O}$ for pixel coordinates from $i$ to $j$. The spatial offset encoding $\mathcal{O}$ serves as the spatial distance term. Instead of designing an explicit radiometric difference term, we use MLP $\mathbb{F}$, to implicitly consider the radiometric difference with image encoding $\mathcal{I}_{i}$ and $\mathcal{I}_{j}$ as input. We believe our formulation can further explore additional information like scene context, which goes beyond the traditional bilateral filter that utilizes an explicitly designed radiometric difference term, although our initial inspiration comes from it.
@@
The paradigm of content-dependent parameter generation has been used in previous depth completion methods and gained robust results. GuideNet [41] and RigNet [49] leverage content-dependent weights to establish a sophisticated multi-modal feature fusion. CSPN [3] and CSPN++ [4] dynamically generate content-dependent affinity matrix to iteratively enhance the initial depth. There are two notable differences between these existing methods and ours. Firstly, the content-dependent parameters are only learned from image content in these methods, typically neglecting the spatial distance between the target pixel and nearby depth measurements. In contrast, our method takes both image content and spatial distance into account. This allows the propagation of depth with the preference for nearest values in both content and spatial domains. Secondly, the content-dependent parameters are typically employed in the multi-modal fusion or refinement stage in existing methods while we extend this paradigm to the earliest pre-processing stage, enabling subsequent stages to overcome the sparsity problem more effectively.
@@ 3.2.3 Prior Encoding
The proposed bilateral propagation module is arranged in a multi-scale scheme, with prior encodings from the corresponding resolution. For the lowest resolution with $s=5$, we directly adopt the image feature $\mathbf{I}^{s}$ as image encoding. Otherwise, to make image encoding representative, we concatenate the multi-modal fused feature $\mathbf{F}^{s+1}$ (introduced in Sec. 3.3) with depth feature $\mathbf{D}^{s+1}$ in scale $s+1$. The depth feature $\mathbf{D}^{s+1}$ is achieved by inverse projecting estimated depth $D^{s+1}$ to camera space. Then we utilize deconvolution operation to upsample the concatenated feature map to scale $s$. Finally, we concatenate the upsampled feature with $\mathbf{I}^{s}$, and adopt an extra convolution operation to produce the image encoding $\mathcal{I}^{s}$.
We inverse project sparse depth map $S^{s}$ to camera space as depth encoding $\mathcal{S}^{s}$. Except $S^{0}=S$, the sparse depth map $S^{s}$ is downsampled from $S$ via a weighted pooling, whose weights are generated from the image encoding $\mathcal{I}^{s}$. Specially, we adopt a periodic shuffling operator to rearrange the image feature map from shape ${\frac{H}{2^{s}}\times\frac{W}{2^{s}}\times 4^{s}}$ to shape $H\times W$ and apply the exponential transformation, that guarantees the generated weight map is positive and has the same resolution as $S$. More implementation details are introduced in the supplementary material. While the points in low-resolution depth maps are denser, we only use the valid ones for depth propagation shown in Eq. 1 and Eq. 2.
We adopt the directed distance from the target pixel to the valid source pixel, i.e. a 2D vector representing offsets along two axes of the image plane, as the spatial distance encoding $\mathcal{O}$. Position encoding, which transforms the position to a high dimensional space via Fourier transform or learned embedding, has been used in some recent popular network structures, e.g. Transformer [45] and NeRF [30]. We find high dimensional position encoding can induce little performance improvement in our formulation, despite the increased computational cost. We attribute this to the simplicity of a spatial distance term being adequate for coefficient generation in our local combination model, akin to bilateral filtering.
@@ 3.3 Multi-Modal Fusion
For each scale, we can get a dense depth map ${{D}^{\prime}}^{s}$ from the bilateral propagation module which is suitable for CNN to process. We adopt image encoding $\mathcal{I}^{s}$ as the image feature, and inverse project ${{D}^{\prime}}^{s}$ to camera space as the depth feature. Then, we follow the early-fusion scheme that simply concatenates the image feature and depth feature, and feeds the concatenated feature as input to a U-Net structure [37] for multi-modal fusion. The adopted U-Net is an encoder-decoder CNN, aggregating local and global features in multi-scale. The encoder part consists of 2 ResNet blocks [11] for feature extraction in each scale and convolution layer with stride 2 to reduce the resolution of the feature map. The decoder part consists of a deconvolution layer with stride 2 to upsample the feature map and skip connection to fuse the upsampled feature and encoded feature with the same resolution by concatenation operation. The U-Net finally yields a fused feature map $\mathbf{F}^{s}$. And a convolution layer is applied on $\mathbf{F}^{s}$ to estimate a residual depth map, which is then added to the previous depth map ${{D}^{\prime}}^{s}$, resulting in a processed depth map ${{D}^{\prime\prime}}^{s}$.
@@ 3.5.2 Training Setting
We implement our method in pytorch [32] and train it on a GPU workstation with 4 Nvidia RTX 3090 GPUs. Our network is mainly stacked with ResNet blocks [11], in which a DropPath [20] is added before residual addition as a regularization for training. We adopt AdamW [26] with 0.05 weight decay as the optimizer, and clip gradient whose $l^{2}$-norm is larger than $0.1$. Our method is trained from scratch in roughly $300K$ iterations with OneCycle learning rate policy [40] gradually reducing the learning rate to $25\%$ of the largest learning rate. We set batch size as $8$ and the largest learning rate as $0.001$ for KITTI dataset. For NYUv2 dataset, whose image resolution is lower, we increase the batch size to $16$ and the largest learning rate to $0.002$ correspondingly. The final model is yielded by Exponential Moving Average (EMA) with $0.9999$ decay.
@@ 4.1 Datasets & Evaluation Metrics
Outdoor scene. The KITTI depth completion (DC) dataset [44] is collected by an autonomous driving vehicle, whose ground truth depth is from temporally registered LiDAR scans, further verified by stereo image pairs. This dataset provides $86,898$ and $1,000$ frames for training and validation respectively. Another $1,000$ frames for testing are evaluated on the remote server with a public leaderboard [^2] http://www.cvlibs.net/datasets/kitti/eval_depth.php?benchmark for ranking. We randomly crop frames to $1216\times 256$ for training and directly use the full-resolution frames as input for testing. The standard evaluation metrics are root mean squared error (RMSE), mean absolute error (MAE), root mean squared error of the inverse depth (iRMSE), and mean absolute error of the inverse depth (iMAE), among which RMSE is chosen as the primary metric for ranking.
@@ 4.2 Comparison with SOTAs
We evaluate our method on the test set of the NYUv2 dataset and KITTI DC dataset separately. Tab. 1 lists the quantitative comparison of our method and other top-ranking published methods. On the KITTI DC leaderboard, our method ranks 1st and exceeds all other methods under the primary RMSE metric at the time of paper submission. It also has comparable performance under other evaluation metrics. On the NYUv2 dataset, our method achieves the best RMSE and the best $\delta_{1.25}$. Due to the diverse viewpoints and sparser input depth map in the NYUv2 dataset, it is more challenging to obtain a large performance gain in the NYUv2 dataset compared with the KITTI dataset.
@@ 4.3.1 Different Number for Neighboring Pixels
In our bilateral propagation module, the target depth is propagated from $N$ nearest neighboring valid depth measurements according to our depth parametrization in Eq. 1. We experiment with different values of $N$ and plot the corresponding RMSEs in Fig. 5. While different choices of $N$ might slightly change the results at the Pre. stage, they can all improve the MF. stage for more effective depth completion. Thus, our method is not very sensitive to this hyperparameter. The performance gain is most noticeable when $N$ increases from 1 to 2, after which it gradually decreases and trends toward saturation. Finally, we choose $N=4$ in our method.
@@
Table 2: Ablation studies on NYUv2 dataset.
4.3.2 Effect of Bilateral Parameters Generation
@@ 4.3.3 Effect of Multi-Scale Depth Estimatation
We also verify the effectiveness of our multi-scale scheme, which estimates dense depth in a multi-scale architecture with a multi-scale loss. We train a variant by replacing the multi-scale loss with a loss only on the original scale and train another variant by further replacing the multi-scale architecture with depth estimation only on the original scale. Based on the results listed in Tab. 2, we can see the integration of both multi-scale architecture and multi-scale loss enhances system performance while merely introducing the multi-scale architecture without incorporating multi-scale loss decreases the performance. In BP-Net, multi-scale architecture facilitates coarse-to-fine depth estimation, where low-resolution results are upsampled to guide the propagation in high-resolution. Consequently, multi-scale loss provides direct supervision at each resolution level, making the low-resolution results meaningful and improving the guidance effectively. Omitting the multi-scale loss might make the network ignore low-resolution results, and increase training difficulty.
@@ 4.3.4 Effect of 3-stage Depth Estimation
At each scale, dense depth is estimated via three sequential stages, consisting of pre-processing (Pre.), multi-modal fusion (MF.), and post-processing (Post.). We train different variants with different combinations of these three stages and analyze the performance gain on RMSE and extra cost on parameters (Params.), Multiply-adds operations (Madds.), GPU RAM and runtime. As shown in Tab. 3, combining MF. with Pre. can achieve $2.12mm$ performance gain on RMSE at the cost of $1.11M$ more Params. and $7.39G$ more Madds. As a comparison, combining MF. with Post. produces less performance gain with more cost on Params. and Madds. In addition, Pre.+MF. achieves better RMSE than MF.+ Post. with less GPU RAM and runtime. And combining all three stages achieves the best RMSE. These comparison verify the effectiveness of our 3-stage design and prove the importance of depth propagation at the preprocessing stage in contrast to the refinement stage.
@@
Table 3: Effect of 3-stage estimation.
4.4 Input Sparsity Analysis
@@ 5 Conclusion
Acknowledgements: This work was supported in part by National Natural Science Foundation of China (NSFC) under Grants 61973311, 62273353, and 62103431.